OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? If g(x) = 2x – 5, then evaluate g(2x2 – 3). a. 2x2 + 8 b. 4x2 – 1 c. 2x2 – 8 d. 4x2 – 11 *not c*

OpenStudy (horsegirl27):

So what do you need help with? what do you have so far?

OpenStudy (horsegirl27):

Because I can help. what do you need to do? solve each problem?

OpenStudy (anonymous):

evaluating

OpenStudy (horsegirl27):

ok. What do you have so far

OpenStudy (anonymous):

First solve this g(2x2-3)

OpenStudy (anonymous):

@ninaesb

OpenStudy (anonymous):

[(2)(2)-3] = ? @ninaesb

OpenStudy (anonymous):

@MARC_ would it be 1 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now,

OpenStudy (anonymous):

substitute g(1) into the equation Note: x=1 g(x)= 2x-5 substitute x=1 @ninaesb

OpenStudy (anonymous):

@MARC_ would it be B ?

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

@MARC_ what about D ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@MARC_ thanks

OpenStudy (anonymous):

Yw @ninaesb

OpenStudy (horsegirl27):

substitute g(1) into the equation Note: x=1 g(x)= 2x-5 Substitute x=1

OpenStudy (horsegirl27):

oh someone already answered. Yeah it was D